the accelaration of a particle moving in a straight line depends on its velocity as a=-v(ln(2^(1/2)). the time interval where velocity becomes half of its value at beginning is?

1sec 2 sec 3 sec or 4 sec

Grade:12th Pass

1 Answers

Vikas TU
7 years ago
Given,
a=-v(ln(2^(1/2))
a = – (2.303v/2)*log2 = > -0.3v
dv/v = – 0.3dt
inetegerating both sides we get,
logv  – log(v/2) = – 0.3t
-log2 = -0.3t
t = 1 sec.